# Computing Floquet quasienergies in finite and extended systems: Role of electromagnetic and quantum-geometric gauges

### v1

@@ Appendix A No net momentum transfer
Here, $t^{\prime}$ is completely arbitrary an we fix it such that $\bm{A}(t^{\prime})=\bm{0}$. Then, $\bm{A}$ in the velocity gauge is related to the functional form of $\bm{E}$ given in Eq. (56) as
@@ Appendix B Expansion of the interaction Hamiltonian in the particle in a box example
In this appendix we consider the expansion of the velocity-gauge interaction Hamiltonian of Eq. (49) up to $M=2$ for the case of the particle in a box studied in Sec. III. Given that the system does not have $\bm{k}$ dependence, the resulting Hamiltonian matrix elements can be written as
@@
Figure 6 compares the quasienergy spectrum calculated using the approximation of Eq. (60) with the exact result of Eq. (34) (corresponding to $V$ curve in Fig. 1). The quasienergies of the two lowest energy states that are close to resonance are perfectly described in the whole range. Other low-lying states, such as the two that display an avoided crossing at $qE_{0}a/\hbar\omega\sim 6$, are also very well described. In contrast, the highest energy states whose quasienergies show barely no dispersion are only well described at low values of $qE_{0}a/\hbar\omega$. As mentioned in the main text, results can be improved in two ways; by considering additional terms in the series expansion of Eq. (49) or employing a larger basis set of states.
Appendix C Numerical discretization of the time-evolution operator
In this appendix we elaborate on the numerical scheme used to calculate the eigenvalues of the effective Floquet Hamiltonian. To this end, we describe the method employed to compute the one period time-evolution operator $\hat{U}(t_{0}+T,t_{0})$, which then allows to obtain $\hat{H}_{F}(t_{0})$ from Eq. (9). The numerical scheme discretizes the time interval $[t_{0},t_{0}+T]$ according to
$t\longrightarrow t_{j}=t_{0}+T\frac{j-1}{N-1}=t_{0}+\delta t(j-1),\quad j\in[1,N],$ (61)
and we then consider the Suzuki-Trotter expansion of the one period time-evolution operator [63],
$\begin{split}\hat{U}(t_{0}+T,t_{0})=&\text{exp}\left[-\frac{i}{\hbar}\sum_{j=1}^{N}\delta t\hat{H}(t_{j})+\mathcal{O}(\delta t^{2})\right]\\ =&\prod_{j=1}^{N}\text{exp}\left[-\frac{i}{\hbar}\delta t\hat{H}(t_{j})\right]+\mathcal{O}(\delta t^{2}),\end{split}$ (62)
@@ Appendix C Numerical discretization of the time-evolution operator
The aim is to calculate the matrix $U(t_{0}+T,t_{0})$ of the time-evolution operator, so this approach requires a matrix representation of $\hat{H}(t_{j})$ for each $j$, which is considered in the main text in the context of the dipole approximation for the length and velocity gauges.
@@ Appendix E Wannier interpolation
$\xi^{j\;\text{W}}(\bm{k})=$ $\sum_{\bm{R}}e^{i\bm{k}\cdot\bm{R}}\xi^{j\;\text{W}}(\bm{R}).$ (74)
and k-space derivatives of these terms can be computed straightforwardly [46, 64, 65]. We have implemented this technique in the WannInt [59] package.
Once these quantities are computed in a certain quantum-geometric gauge, we have means to resolve $\hat{H}(t)$. In the velocity-gauge expression, each individual term of the expansion Eq. (49) is explicitly quantum-geometric gauge-covariant. Therefore, we have proceeded by using the Wannier interpolation technique to compute the matrices representing $\hat{\mathcal{D}}^{\alpha_{1}}\cdots\hat{\mathcal{D}}^{\alpha_{n}}[\hat{H}_{0}]$ in the Wannier gauge. Then, we obtain the matrices of $H^{\text{W}}(\bm{k},t)$, $U^{\text{W}}(\bm{k},t_{0}+T,t_{0})$, and $H_{\text{F}}^{\text{W}}(\bm{k},t_{0})$ in the Wannier gauge by using Eqs. (49), (62) and  (9), respectively. For the computation of the expansion terms of Eq. (49), we have used
@@
and the analogous expressions for higher order terms $(\hat{\mathcal{D}}^{l}\hat{\mathcal{D}}^{j}[\hat{H}_{0}])^{\text{W}}_{nm},\;\cdots$ by employing Eq. (47). In the case of the length-gauge approximation of Eq. (48), $\hat{H}(t)$ is not gauge-covariant and is evaluated in the quantum-geometric gauge that diagonalizes the Hamiltonian. In this case, we first calculate the matrices $H_{0}^{\text{W}}(\bm{k})$ and $\xi^{j\;\text{W}}(\bm{k})$ in the Wannier gauge and transform them to a new gauge by diagonalizing $H_{0}^{\text{W}}(\bm{k})$ and using their respective gauge transformation properties. That is, for this case we define the gauge transformation $\mathcal{U}(\bm{k})$ in Eq. (42) as the gauge transformation diagonalizing $H_{0}^{\text{W}}(\bm{k})$. This defines the so called “Hamiltonian” gauge H, where the matrices $H_{0}^{\text{H}}(\bm{k})$ and $\xi^{j\;\text{H}}(\bm{k})$ are calculated. It is in this gauge that we compute the approximation Eq. (48) and calculate, using the length gauge, the matrix $H^{\text{H}}(\bm{k},t)$ in the quantum-geometric Hamiltonian gauge. We then proceed as in the velocity gauge case and calculate the matrix $H^{\text{H}}_{\text{F}}(\bm{k})$, which as the superscript indicates is calculated in the Hamiltonian gauge.

### v2

@@ Appendix A No net momentum transfer
Here, $t^{\prime}$ is completely arbitrary and we fix it such that $\bm{A}(t^{\prime})=\bm{0}$. Then, $\bm{A}$ in the velocity gauge is related to the functional form of $\bm{E}$ given in Eq. (56) as
@@ Appendix B Expansion of the interaction Hamiltonian in the particle in a box example
In this appendix we consider the expansion of the velocity gauge interaction Hamiltonian of Eq. (49) up to $M=2$ for the case of the particle in a box studied in Sec. III. Given that the system does not have $\bm{k}$ dependence, the resulting Hamiltonian matrix elements can be written as
@@
Figure 6 compares the quasienergy spectrum calculated using the approximation of Eq. (60) with the exact result of Eq. (34) (corresponding to $V$ curve in Fig. 1). The quasienergies of the two lowest energy states that are close to resonance are perfectly described in the whole range. Other low-lying states, such as the two that display an avoided crossing at $-qE_{0}a/\hbar\omega\sim 6$, are also very well described. In contrast, the highest energy states whose quasienergies show barely no dispersion are only well described at low values of $-qE_{0}a/\hbar\omega$. As mentioned in the main text, results can be improved in two ways; by considering additional terms in the series expansion of Eq. (49) or employing a larger basis set of states.
Appendix C Numerical discretization of the time evolution operator
In this appendix we elaborate on the numerical scheme used to calculate the eigenvalues of the effective Floquet Hamiltonian. To this end, we describe the method employed to compute the one period time evolution operator $\hat{U}(t_{\text{s}}+T,t_{\text{s}})$, which then allows to obtain $\hat{H}_{F}(t_{\text{s}})$ from Eq. (9). The numerical scheme discretizes the time interval $[t_{\text{s}},t_{\text{s}}+T]$ according to
$t\longrightarrow t_{j}=t_{\text{s}}+T\frac{j-1}{N_{t}-1}=t_{\text{s}}+\delta t(j-1),\quad j\in[1,N_{t}],$ (61)
and we then consider the Suzuki-Trotter expansion of the one period time evolution operator [61],
$\begin{split}\hat{U}(t_{\text{s}}+T,t_{\text{s}})=&\text{exp}\left[-\frac{i}{\hbar}\sum_{j=1}^{N_{t}}\delta t\hat{H}(t_{j})+\mathcal{O}(\delta t^{2})\right]\\ =&\prod_{j=1}^{N_{t}}\text{exp}\left[-\frac{i}{\hbar}\delta t\hat{H}(t_{j})\right]+\mathcal{O}(\delta t^{2}),\end{split}$ (62)
@@ Appendix C Numerical discretization of the time evolution operator
The aim is to calculate the matrix $U(t_{\text{s}}+T,t_{\text{s}})$ of the time evolution operator, so this approach requires a matrix representation of $\hat{H}(t_{j})$ for each $j$, which is considered in the main text in the context of the dipole approximation for the length and velocity gauges.
@@ Appendix E Wannier interpolation
$\xi^{j\;\text{W}}(\bm{k})=$ $\sum_{\bm{R}}e^{i\bm{k}\cdot\bm{R}}\xi^{j\;\text{W}}(\bm{R}),$ (74)
and k-space derivatives of these terms can be computed straightforwardly [46, 62, 63].
Once these quantities are computed in a certain quantum-geometric gauge, we have means to resolve $\hat{H}(t)$. In the velocity gauge expression, each individual term of the expansion Eq. (49) is explicitly quantum-geometric gauge-covariant. Therefore, we have proceeded by using the Wannier interpolation technique to compute the matrices representing $\hat{\mathcal{D}}^{\alpha_{1}}\cdots\hat{\mathcal{D}}^{\alpha_{n}}[\hat{H}_{0}]$ in the Wannier gauge. Then, we obtain the matrices of $H^{\text{W}}(\bm{k},t)$, $U^{\text{W}}(\bm{k},t_{\text{s}}+T,t_{\text{s}})$, and $H_{\text{F}}^{\text{W}}(\bm{k},t_{\text{s}})$ in the Wannier gauge by using Eqs. (49), (62) and  (9), respectively. For the computation of the expansion terms of Eq. (49), we have used
@@
and the analogous expressions for higher order terms $(\hat{\mathcal{D}}^{l}\hat{\mathcal{D}}^{j}[\hat{H}_{0}])^{\text{W}}_{nm},\;\cdots$ by employing Eq. (47). In the case of the length gauge approximation of Eq. (48), $\hat{H}(t)$ is not gauge-covariant and is evaluated in the quantum-geometric gauge that diagonalizes the Hamiltonian. In this case, we first calculate the matrices $H_{0}^{\text{W}}(\bm{k})$ and $\xi^{j\;\text{W}}(\bm{k})$ in the Wannier gauge and transform them to a new gauge by diagonalizing $H_{0}^{\text{W}}(\bm{k})$ and using their respective gauge transformation properties. That is, for this case we define the gauge transformation $\mathcal{U}(\bm{k})$ in Eq. (42) as the gauge transformation diagonalizing $H_{0}^{\text{W}}(\bm{k})$. This defines the so called “Hamiltonian” gauge H, where the matrices $H_{0}^{\text{H}}(\bm{k})$ and $\xi^{j\;\text{H}}(\bm{k})$ are calculated. It is in this gauge that we compute the approximation Eq. (48) and calculate, using the length gauge, the matrix $H^{\text{H}}(\bm{k},t)$ in the quantum-geometric Hamiltonian gauge. We then proceed as in the velocity gauge case and calculate the matrix $H^{\text{H}}_{\text{F}}(\bm{k})$, which as the superscript indicates is calculated in the Hamiltonian gauge.
Appendix F Time complexity analysis
In this appendix we briefly analyze the time complexity related to the calculation of the quasienergy spectrum. We start by analyzing the complexity of the method in terms of the internally fixed variables, namely those variables which do not correspond to the physical setting of the calculation and just control the precision. These include the number of states of the crystal spectrum, $N$, leading to $N\times N$ matrices representing $\hat{H}_{0}$ and $\hat{\xi}^{j}$. The number of steps for the discretization of the one-cycle time evolution operator in Eq. (62), $N_{t}$, and the calculation method employed to resolve $\hat{H}(t_{j})$, where we distinguish between finite and extended systems. In the case of extended systems we employ Eq. (49) (ExV) and in the case of finite systems we further distinguish between velocity and length gauge expressions, Eqs. (30) (FnV) and (28) (FnE), respectively.
By virtue of Eq. (62), the computation of the one-cycle time evolution operator requires computing $N_{t}$ matrix exponentials of size $N\times N$, which are then multiplied sequentially. Then, by virtue of Eq. (9), a further diagonalization is required to obtain the quasienergy spectrum. Each $N\times N$ matrix multiplication has, at most, a time complexity of $\mathcal{O}(N^{3})$ [64], while the diagonalization procedure, carried out by Lapack’s Zheev routine, which employs the QR algorithm, involves a time complexity of $\mathcal{O}(N^{3})$ [64]. The matrix exponential operation involves two matrix multiplications and a single diagonalization. This leads to a complexity
$\begin{split}&\mathcal{O}(C)+\mathcal{O}(N_{t}(N^{3}+2N^{3}+N^{3})+N^{3})=\\ &\mathcal{O}(C)+\mathcal{O}((4N_{t}+1)N^{3}),\end{split}$ (76)
where $C$ is the time complexity associated with the resolution of $\hat{H}(t_{j})$ in any of the FnE, FnV and ExV cases.
The FnE case requires adding $d+1$ matrices, thus
$C^{\text{FnE}}=\mathcal{O}(dN^{2}),$ (77)
the FnV case requires computing two matrix exponentials and two matrix multiplications, leading to
$C^{\text{FnV}}=\mathcal{O}(2\times 3N^{3}+2N^{3})=\mathcal{O}(8N^{3}).$ (78)
Lastly, the ExV case requires computing $M$ objects, the successive actions $\hat{D}^{\alpha_{M}}\cdots\hat{D}^{\alpha_{1}}[\hat{H}_{0}]$, each of which can be considered an $N\times N\times d^{l}$ array, where $l\in[1,M]$. Each object is contracted $d\times l$ times with the $A_{\alpha_{j}}(t)$s, the vector potential components, obtaining $M$ matrices of size $N\times N$, which are then summed together. The calculation of a single $\hat{D}^{\alpha_{l}}\cdots\hat{D}^{\alpha_{1}}[\hat{H}_{0}]$ array involves summing $l$-times nested commutators, thus it involves a computational complexity of $\mathcal{O}(l\times N^{3})$. The contraction with the $A_{\alpha_{j}}(t)$s involves a computational complexity of $\mathcal{O}(d\times l\times N^{2})$ and the addition of $M$ matrices involves a computational complexity of $\mathcal{O}(M\times N^{2})$. Thus, ignoring $\mathcal{O}(N^{2})$ contributions,
$\begin{split}&C^{\text{ExV}}=\mathcal{O}\left(N^{3}\times\sum_{l=1}^{M}l\right)=\\ &\mathcal{O}\left(\frac{(M+1)M}{2}N^{3}\right).\end{split}$ (79)
Finally, we deal with the functional dependence that the quasienergy spectrum has with respect to externally fixed variables, namely with respect to the amplitudes $E_{n}^{i}$, phases $\varphi_{n}^{i}$ and frequency $\omega$ in Eq. (56). This accounts for $12\times N_{h}\times d+$ 1 variables, where $N_{h}$ stands for the number of harmonics considered in Eq. (56) and $d$ for the dimensionality. Finally, the spectrum also depends functionally on $\bm{k}$ in the case of extended systems, thus
$\varepsilon_{n}\equiv\varepsilon_{n}\left(E_{1}^{1},\cdots,E_{N_{h}}^{d},\varphi_{1}^{1},\cdots,\varphi_{N_{h}}^{d},\omega,\bm{k}\right).$ (80)
Each variation of any of the external variables involves a multiplicative increase of time complexity, since the algorithm must be run again for each external variable permutation.